Expand and Simplify he expression: Show all work for credit. In In dx-In & In x - In y in 01/10 - 1/2 iny
Answer #1

chus en sny ln (my) 1/2 N + ln (ry) = 1/2 (lun-lny) as in my = I lan- Izliny So, fourth of hen is correct